Solve with Substitution, explain how you got your answer with each step.
7x + y = 7
4x + 3y = 21

Answer by rothauserc(4718)

1) 7x + y = 7
:
2) 4x + 3y = 21
:
solve equation 1) for y
:
y = 7 - 7x
:
substitute for y in equation 2
:
4x +3(7 - 7x) = 21
:
4x +21 - 21x = 21
:
-17x = 0
:
x = 0
y = 7 -7(0) = 7
:
************************
x = 0 and y = 7
************************
: